Project is required in initial investment of ruppes 500000 cash flow after for its estimated life of 4 years are as follows

Year. CFAT
1. 100000
2. 200000
3. 150000
4. 160000
1
Expert's answer
2019-11-18T13:12:08-0500

We can calculate net present value (NPV) using this data, but also should know the discount rate during these 4 years, because the formula for NPV is:

"NPV = -I + \frac{CF1} {(1 + r)^1} + \frac{CF2} {(1 + r)^2} + ... +\frac{CFn} {(1 + r)^n},"

where r is discount rate, I is initial investment, CF is cash flow.